in regression analysis, the Sum of Products (SP) value will determine whether the slope will be positive or negative
Answer #1

Okay, so when we consider a regression analysis, the slope of the graph (r) between the two variables (say X and Y) is actually defined as a ratio of the sum of products and the square root of the product of the sum of squares (of X) and the sum of squares (of Y). Thus, since the sum of squares is always positive, the Sum of Products determines whether the slope will be positive or negative.

.SSx= (X – X)(X – X) SS = (Y -Ỹ)(Y -Ỹ) SPxy = S(x - 7)(-) T=SPry/V(SSr.SSy)